Question 10 (PLEASE EXPLAIN EACH ANSWER STEP BY STEP)

Susan has a part-time business producing seasonal plywood yard ornaments for resale at local craft fairs and bazaars. She currently works 8 hours per day to produce 20 ornaments.

a. What is her productivity?

b. She thinks that by redesigning the ornaments and switching from use of a wood glue to a hot-glue gun she can increase her total production to 30 ornaments per day. What is her new productivity?

c. What is her percentage increase in productivity?
Answer #1

Answer a ) Currently working hour = 8

number of ornaments she produced per day = 20

Productivity P1 = (No. of ornament per day/Currently working hour)

= (20/8)

Productivity P1=  2.5 ornament/hour.

Answer b) Current working hour = 8

No. of ornament after using hot glue gun = 30

Productivity P2 = (No. of ornaments by hot glue gun/Currently working hour)

= (30/8)

Productivity P3 = 3.75 ornament/hour.

Answer c) Productivity without using Hot glue gun P1 = 2.5 Ornament/day

Productivity with using Hot glue gun P2 = 3.75 Ornament/day

Increase in percentage = ((change in productivity)/2 *100)

= ((3.75-2.5)/2 * 100)

= (0.625 *100)

Percent increase in productivity = 62.5%